# Information Discovery in e-Commerce

### v1

@@ 5.3.4 Hybrid matching
Recently, more and more hybrid matching models have been proposed to combine the strengths of representation- and interaction-based models. mitra2017learning proposed a matching model, namely DUET, which integrates both local (interaction-based) and distributed (representation-based) features to calculate query-document relevance. Pre-trained language models, such as BERT [kenton2019bert], have further advanced hybrid approaches by capturing both contextualized token-level interactions and global semantic representations, achieving promising performance in tasks like search and recommendation [sun2019bert4rec, nogueira2019multi, lin2021pretrained]. tracz2020bert proposed a BERT-based model to leverage both types of matching in a similarity learning framework for product matching in e-commerce. yao2022reprbert explored the deployment of BERT in online retrieval systems by distilling it into a representation-based architecture, while still maintaining the advantages of interaction-based processing for more precise matching.
@@ 5.3.5 Matching in personalized search
One of the primary characteristics of e-commerce search is its highly personal variance in queries. First, multiple items could be topic-related with a consumer’s query, but only a few are actually purchased, i.e., different individuals have different opinions even on the same product. Hence e-commerce search without personalization cannot satisfy consumers. Second, personalization has explicit benefits for e-commerce platforms by exhibiting the products that consumers would like to purchase. As we have demonstrated in Section 5.1, the definition of “relevance” for e-commerce search is not the same as for web-based search as most of e-commerce platforms apply gross merchandise volume (GMV) as the gold standard for measuring success. To the best of our knowledge, jannach2017investigating were the first to attempt to personalize product search by using personalized recommendation approaches. However, the matching problem in personalized e-commerce search is more challenging as most platforms have approaches to matching products to queries that are far from perfect. To address this problem, an increasing number of matching studies have been proposed. Matching approaches in personalized product search can be classified into query-independent and query-dependent ones [liu2022category].
Query-independent matching. The query-independent matching methods embedded users into a general profiling vector in the offline training stage [ai2017learning, ai2019explainable, liu2020structural]. ai2017learning proposed a deep neural network and jointly learn latent representations for queries, products, and users. Specifically, a hierarchical embedding model is proposed for personalized e-commerce search. As illustrated in Fig. 5.3, the authors projected both queries and consumers into a single latent space and explicitly control their weights in a personalized product search model. Following van2016learning, the authors designed the latent representations of queries and users to have good compositionality so that the personalized search model can be directly computed as a linear combination of query models and user models. Specifically, both queries and users are projected into a single latent space. Given a query $q$, the corresponding query intent is represented in $R^{\alpha}$; similarly, the user preference is represented in $R^{\alpha}$ given a user $u$. As shown in Fig. 5.3, the personalized product retrieval model is defined as ${M_{uq}}=\lambda q+(1-\lambda)u$, where $\lambda$ is a hyper-parameter that controls the weight of the query model $q$ and the user model $u$. To alleviate the mismatch problem during personalized product search, the authors proposed a hierarchical embedding approach to reflect the distributed representations of users, items, and queries. The experiments are conducted with synthetic queries generated from product category information.
@@
External structured knowledge graphs have been applied to enhance the personalized matching procedure. Structured relationship among users, products, and queries has been jointly considered in a novel graph neural network approach [ai2019explainable, liu2020structural]. ai2019explainable proposed a unified knowledge graph on multiple types of product data, and conducted retrieval with it: a dynamic relation embedding module was proposed to construct a session-based knowledge graph; a soft matching algorithm was able to efficiently extract explainable paths with knowledge embeddings. liu2020structural exploited the structured representation learning scheme from user-query-product interactions with conjunctive graph patterns. Geometric operations, such as projections and intersections, are applied in the proposed graph neural networks. Derived from the knowledge graph embedding, liu2022category used multiple vectors to encode the diverse preferences of users. The authors leveraged the category information to aggregate the multiple interests of users with category indications as references. To sufficiently exploit collaborative signals among products, users, and queries, cheng2022ihgnn proposed a hypergraph-based method from the ternary user-product-query interactions, by considering high-order features of neighbors. Query-independent matching models can calculate user embeddings and store them in advance, which makes it convenient and efficient to apply in real-world search engines. To tackle inconsistent user behavior in multi-stage e-commerce search systems, wang2023learning employed external information to refine query-item matching. By mining various user interactions (ordered, clicked, unclicked items) within a post-fusion strategy, they generated more accurate semantic representations. This approach not only enhances retrieval efficiency, but also improves both offline recall and online conversion rates.
@@
To tackle this challenge, ai2019zero proposed a zero-attention neural network model (ZAM) for personalized product search by conducting differentiated personalization for different query-user pairs. This neural network is designed based on an embedding-based generative framework. In detail, the authors conduct query-dependent personalization by constructing user profiles with a zero attention strategy that enables it to automatically decide when and how to attend in different search scenarios. Likewise, guo2018attentive designed a dual attention-based network to capture users’ current search intentions and their long-term preferences. bi2019study studied relevance feedback based on both long-term and short-term context dependencies in multi-page product search with an end-to-end personalized product search model. xiao2019dynamic proposed a streaming Bayesian method to explicitly and collaboratively learn representations of different categories of entities in a joint metric space over time.
@@
Spareness is another critical challenge in tracking users’ sequential behaviors in product search. Graph-based methods were proposed to tackle this challenge. By utilizing short-term user behaviors, fan2022modeling extended the structural relationship representation learning scheme [liu2020structural] to explore both local and global user behavior patterns on a user successive behavior graph. zhu2022cross integrated cross-domain transfer learning with a knowledge graph to establish the underlying interest correlation. Their proposed method performs interest alignment across domains by explicitly modeling the long-term and short-term interactions between users and items, which capture the dynamics of product properties and user interests.
@@
Pre-trained language models have also been applied to query-dependent matching in personalized product search. To conduct differential personalization adaptively under different contexts, bi2020transformer proposed a transformer-based embedding model (TEM) that is more flexible; in TEM, personalization can vary from no to full effect. In contrast with ZAM, TEM takes into consideration the interactions between purchased items so that it learns better dynamic representations of queries and items, which leads to better attention weights in personalized product search. Based on TEM, the review-based transformer model, abbreviated as RTM, is proposed to match user intents and items at the level of finer-grained information (e.g., their associated reviews) [bi2020learning]. The authors conducted user-item matching at the review level so that the reason an item is ranked at the top can be explained; also, the importance of each user and item review during matching is dynamically adapted. RTM represents users and items based only on their reviews without the need for their identifiers, which improves the generalization ability during product search. dai2023contrastive proposed a contrastive learning framework, namely CoPPS, to enhance user representations for personalized product search. By pre-training a sequence encoder with contrastive sampling and fine-tuning, CoPPS employs multiple data augmentation strategies to improve user modeling.
@@
For real-world applications, zhang2020towards jointly considered semantic matching and personalized matching at JD.com. The authors propose the deep personalized and semantic retrieval (abbreviated as DPSR) model with a two-tower architecture: a multi-head design of a query tower and an attention-based loss function. Similarly, magnani2022semantic proposed a semantic retrieval model with a two-tower architecture in e-commerce search production at Walmart.com. The authors proposed to select negative examples for training a large semantic retrieval model and an approximate metric to evaluate the performance.
@@ 5.4 Ranking strategies in e-commerce search
As we have discussed in Section 5.1.1, learning to rank approaches have been applied to e-commerce search. Unlike web search scenarios based on relevance judgments, e-commerce search has multiple implicit and explicit signals that need to be integrated during the ranking process. In e-commerce search, learning to rank methods need to tackle a series of challenges [karmaker2017application]. Learning to rank methods assume that the scoring of items to be ranked is a parameterized function of multiple features computed based on the given query and the items, whereas parameters are used to control the weights of features. A large number of product and query features, such as brands, rating, categories, etc., are important to present useful representations in e-commerce learning to rank models. Also, product popularity related features have been verified effective in optimizing ranking results of e-commerce search. karmaker2017application studied performances of multiple learning to rank strategies in e-commerce search, and found that LambdaMART [burges2010ranknet] is able to learn a balance between these two kinds of features. Detailed comparisons are listed in Table 5.1. In addition, user engagement behavior, such as clicks and orders, also play an important role in optimizing ranking results [karmaker2017application, wu2018turning]. In contrast with web search which only has clicks to judge the relevance, e-commerce search contains $4$ prominent relevance feedback behavior: clicks, cart-adds, orders and revenue. Thus, various training objectives can be considered in optimizing e-commerce ranking results.
@@
Another main challenge for e-commerce learning to rank methods is the lack of labeled information. In web search, high-quality labeled information obtained by eliciting relevance ratings from human experts or crowdsourcing, makes learning to rank methods effective. However, in the context of e-commerce search it is infeasible to found a standard method to obtain ground-truth information. Intuitively, crowd sourcing annotations seem to provide labels for e-commerce learning to rank methods. karmaker2017application and alonso2009relevance studied the reliability of the relevance judgements provided by the crowd workers for e-commerce queries. However, the authors found that crowdsourcing fails to provide reliable relevance judgements for e-commerce queries. Thus implicit feedback from e-commerce users is an important source of information that helps reveal the saliency of products for a given query.
@@
wu2018turning divided the ranking optimization problem into two successive stages: click optimization and purchase optimization. We have discussed previous studies on click modeling and purchase modeling in Section 4.1.1 and Section 4.1.3, respectively. Inspired by these approaches, wu2018turning applied a list-wise learning to rank model [cao2007learning] to optimize clicks by jointly considering item positions and query-level structures; a binary classification approach was applied to predict the purchase behavior.
@@
Online learning to rank approaches have been applied to optimize ranking results in e-commerce search [hu2018reinforcement]. Unlike traditional static learning to rank models, online learning to rank methods optimize the production ranker interactively by exploiting user implicit feedbacks [zoghi-copeland-2015, oosterhuis-balancing-2017, schuth-multileave-2016]. Online learning to rank approaches can be divided into two groups: the first is to learn the best ranking function from a function space [hofmann2016online, hofmann2013balancing, yue2009interactively]; the second group directly learns the best list under some model of user interactions [radlinski2008learning, schuth-multileave-2016, oosterhuis-balancing-2017]. As part of the first group of online learning to rank methods, hu2018reinforcement proposed a reinforcement learning method for ranking optimization in e-commerce searching scenarios. The authors formulated the multi-step ranking procedure in e-commerce search as a search session Markov decision process (SSMDP). An algorithm, named deterministic policy gradient with full backup estimation (DPG-FBE), was then proposed for the problem of high reward variance and unbalanced reward distribution of SSMDP. To integrate search results from heterogeneous sources, takanobu2019aggregating proposed a search result aggregation method that formulates a semi-Markov decision process, where a low-level policy is applied to represent items and a high-level policy is used to select rankers. Based on a large-scale real-world dataset, pkddAnwaarRK20 employed a counterfactual risk minimization (CRM) approach to directly optimize the ranking list from the log data.
@@
In recent years, deep neural-based retrieval models [mitra2017introduction, kenter-neural-2017, onal-neural-2018] were used in the ranking step of the e-commerce search [magnani2019neural, zhang2019neural, wu2022multi]. Specifically, magnani2019neural enhanced the deep neural network model using different types of text representation and loss function at Warmart.com. Whereas zhang2019neural proposed an e-commerce ranking model with interaction features between the query and a graph of products maintaining product interactions. Pre-trained foundation models have been verified effective in real-world web search scenarios [zou2021pre, lin2021pretrained, chu2022h]. In e-commerce search, wu2022multi applied BERT for product ranking within a multi-task learning framework. The authors used the probability transfer method in the framework to model multiple sequential engagement behaviors. By integrating semantic matching features output by the domain-specific BERT, the authors confirmed the effectiveness of their proposed approach on real-world e-commerce search data.
@@ 5.5.1 Multi-modal e-commerce search
Along with the rise of deep neural networks, recently multi-modal search has received an increasing number of attention [zhang2013attribute, mao2014deep, yang2017visual, liu2018attentive, balaneshin2018deep, guo2018multi, zhang2018visual, qu2021dynamic, wei2021universal, tan2022bit]. The key of multi-modal search is to find an effective mapping mechanism to project data from different modalities into a common latent space. Multi-modal search approaches can be classified into hashing-aware models and semantic-aware models. The former type of methods map various modalities in the original space to a Hamming space using hash functions [cao2017transitive, luo2018scalable, zhu2020deep, tan2022bit]; whereas the latter ones project the multi-modal data into a low-dimensional space by learning a mapping function [wang2016effective, laenen2018web, qu2021dynamic, wang2022siamese].
@@
As we discussed in previous sections, most approaches to e-commerce search focus on the textual matching problem. However, multi-modal search in e-commerce is becoming more and more important along with the rise in online photos and openly available image datasets [yang2017visual, zhang2018visual, li2018design, wang2020metasearch, sigir21DaganGN21, liu2022pretraining]. Real-world scenarios have received a lot of attention: yang2017visual proposed a novel end-to-end approach for scalable visual search infrastructure at Ebay. Similar platforms can be found at Alibaba [zhang2018visual] and JD.com [li2018design, wang2020metasearch]. Early studies on multi-modal e-commerce search has mostly focused on the fashion category [yang2017visual, zhang2018visual, li2018design]. However, wang2020metasearch, sigir21DaganGN21, liu2022pretraining confirmed that multi-modal search is widespread across many e-commerce categories, especially categories that involve aspects that are harder to express verbally, but can naturally be captured visually, such as style, type, and pattern.
@@
In multi-modal search, multiple modalities, such as text and images, can be found in both queries and products. Most of existing multi-modal search solutions optimize a function to project the multi-modal data into a low-dimensional space after unified representation learning. laenen2018web presented a multi-modal search paradigm for e-commerce search that results in an improved shopping experience. The authors reason with both images and languages through a common embedding space. guo2018multi formulated the e-commerce personalized search problem based on the relevance between images and text with respect to the query and the user preferences from both textual and visual modalities. The authors modeled the relevance of a product with respect to the query and the user’s multi-modal preferences. A transition-based product search method was proposed, where the multi-modal feature space is initialized based on the textual and visual features of products. An interpretable multi-modal e-commerce retrieval framework has been proposed for fashion products [liao2018interpretable]. The authors bridged the gap between deep features and meaningful fashion concepts. They proposed a hierarchical similarity function to accurately characterize the semantic affinities among fashion items.
@@
sigir21DaganGN21 highlight various differences between visual and textual search, which can be summarized in the following main challenges in multi-modal e-commerce search:
@@
  • Large-scale sparse data. Large multimedia corpora make scalability and efficiency key requirements for e-commerce search. Visual queries are more specific than text queries, which results in a lower number of retrieved results and sparse coverages of categories.
  • Limited user engagement. In multi-modal e-commerce search, user engagement behaviors, e.g., clicks, dwell time, purchases, etc., is substantially sparser than in textual search.
  • Ambiguous user intents. Two different main use cases exist in multi-modal e-commerce search: target finding and decision making [sigir21DaganGN21]. The two use cases reflect the navigational intent and informational intent respectively.
Additionally, annotating new product images and retraining a new feature representation model on large-scale data is expensive. To address this problem, few-shot multi-modal product search was proposed to update the feature representation and index model with few-shot data and employ a fast learning strategy for new categories. wang2020metasearch proposed a framework for few-shot incremental search via meta-learning, with a multi-pooling feature extractor to extract discriminative multi-modal features. Recently, based on pre-trained language models, liu2022pretraining proposed an effective contrastive learning framework to learn representations of multi-modal search sessions based on multi-view heterogeneous graph networks. liu2023multimodal employed BERT with self-distillation framework for product understanding by integrating visual and textual information.
Cross-modal search is receiving more and more attention [qu2021dynamic, wang2022siamese, tan2022bit]. Unlike multi-modal search, cross-modal search aims to solve the discrepancy problem between different modalities in search sessions [wang2022siamese]. Although much progress has been made in bridging multiple modalities, it still remains challenging because of the difficult intra-modal reasoning and cross-modal alignment [qu2021dynamic]. As complicated relations exist among products in e-commerce, it is more difficult to explore these multi-hop interactions in cross-modal product search scenarios. Moreover, the extremely high computation cost brought by multi-modal input limits its usefulness in large-scale cross-modal retrieval in e-commerce applications.

### v3

@@ 5.3.4 Hybrid matching
Several hybrid matching models have been proposed to combine the strengths of representation- and interaction-based models. mitra2017learning propose a matching model, DUET, that integrates both local (interaction-based) and distributed (representation-based) features to calculate query-document relevance. Pre-trained language models, such as BERT [kenton2019bert], have further advanced hybrid approaches by capturing both contextualized token-level interactions and global semantic representations, achieving promising performance in tasks like search and recommendation [sun2019bert4rec, nogueira2019multi, lin-2021-pretrained]. tracz2020bert introduce a BERT-based model to use both types of matching in a similarity learning framework for product matching in e-commerce. yao2022reprbert explore the deployment of BERT in online retrieval systems by distilling it into a representation-based architecture, while still maintaining the advantages of interaction-based processing for more precise matching.
@@ 5.3.5 Matching in personalized search
One of the primary characteristics of e-commerce search is its highly personal variance in queries. First, multiple items could be topic-related with a consumer’s query, but only a few are actually purchased, i.e., different individuals have different opinions even on the same product. Hence e-commerce search without personalization is unlikely to satisfy consumers. Second, personalization has explicit benefits for e-commerce platforms by exhibiting the products that consumers would like to purchase. As we have seen in Section 5.1, the definition of “relevance” for e-commerce search is not the same as for web-based search as most e-commerce platforms apply gross merchandise volume (GMV) as the gold standard for measuring success. To the best of our knowledge, jannach2017investigating have been the first to attempt to personalize product search by using personalized recommendation approaches. However, the matching problem in personalized e-commerce search is more challenging as most platforms have approaches to matching products to queries that are far from perfect. To address this problem, an increasing number of matching studies have been proposed. Matching approaches in personalized product search can be classified into query-independent and query-dependent ones [liu2022category].
Query-independent matching. Query-independent matching methods embed users into a general profiling vector in the offline training stage [ai2017learning, ai2019explainable, liu2020structural]. ai2017learning design a deep neural network and jointly learn latent representations for queries, products, and users. A hierarchical embedding model is proposed for personalized e-commerce search. As illustrated in Fig. 5.3, the authors project both queries and consumers into a single latent space and explicitly control their weights in a personalized product search model. Following van2016learning, the authors design latent representations of queries and users to have good compositionality so that the personalized search model can be directly computed as a linear combination of query models and user models. Both queries and users are projected into a single latent space. Given a query $q$, the corresponding query intent is represented in $R^{\alpha}$; similarly, the user preference is represented in $R^{\alpha}$ given a user $u$. As shown in Fig. 5.3, the personalized product retrieval model is defined as ${M_{uq}}=\lambda q+(1-\lambda)u$, where $\lambda$ is a hyper-parameter that controls the weight of the query model $q$ and the user model $u$. To alleviate the mismatch problem during personalized product search, the authors put forward a hierarchical embedding approach to reflect the distributed representations of users, items, and queries. Their experiments are conducted with synthetic queries generated from product category information.
@@
External structured knowledge graphs have been applied to enhance the personalized matching procedure. Structured relationships among users, products, and queries have been jointly considered in graph neural network approaches. ai2019explainable introduce a unified knowledge graph on multiple types of product data and conduct retrieval with it. A dynamic relation embedding module constructs a session-based knowledge graph and a soft matching algorithm extracts explainable paths with knowledge embeddings. liu2020structural exploit the structured representation learning scheme from user-query-product interactions with conjunctive graph patterns. Geometric operations, such as projections and intersections, are applied in the proposed graph neural networks. Derived from knowledge graph embeddings, liu2022category use multiple vectors to encode the diverse preferences of users. The authors used the category information to aggregate the multiple interests of users with category indications as references. To exploit collaborative signals among products, users, and queries, cheng2022ihgnn offer a hypergraph-based method from the ternary user-product-query interactions, by considering high-order features of neighbors. Query-independent matching models can calculate user embeddings and store them in advance, which makes it convenient and efficient to apply in real-world search engines. To tackle inconsistent user behavior in multi-stage e-commerce search systems, wang2023learning employ external information to refine query-item matching. By mining various user interactions (ordered, clicked, unclicked items) within a post-fusion strategy, they generate more accurate semantic representations. This approach not only enhances retrieval efficiency, but also improves both offline recall and online conversion rates.
@@
To tackle this challenge, ai2019zero introduce a zero-attention neural network model (ZAM) for personalized product search that conducts differentiated personalization for different query-user pairs. This network builds on an embedding-based generative framework. Query-dependent personalization is implemented by constructing user profiles with a zero attention strategy that enables it to automatically decide when and how to attend in different search scenarios. Likewise, guo2018attentive design a dual attention-based network to capture users’ current search intentions and their long-term preferences. bi2019study study relevance feedback based on both long-term and short-term context dependencies in multi-page product search with an end-to-end personalized product search model. xiao2019dynamic propose a streaming Bayesian method to explicitly and collaboratively learn representations of different categories of entities in a joint metric space over time.
@@
Sparseness is another critical challenge in tracking users’ sequential behaviors in product search. Graph-based methods haven been proposed to tackle this challenge. By using short-term user behavior, fan2022modeling extend a structural relationship representation learning scheme [liu2020structural] to explore both local and global user behavior patterns. zhu2022cross integrate cross-domain transfer learning with a knowledge graph to establish the underlying interest correlation. Their proposed method performs interest alignment across domains by explicitly modeling the long-term and short-term interactions between users and items, which capture the dynamics of product properties and user interests.
@@
Pre-trained language models have also been applied to query-dependent matching in personalized product search. To conduct differential personalization in different contexts, bi2020transformer propose a transformer-based embedding model (TEM); in TEM, personalization can vary from no to full effect. In contrast with ZAM, TEM takes into consideration the interactions between purchased items so that it learns better dynamic representations of queries and items, which leads to better attention weights in personalized product search. Based on TEM, a review-based transformer model, abbreviated as RTM, is designed to match user intents and items at the level of finer-grained information (e.g., their associated reviews) [bi2020learning]. The authors conduct user-item matching at the review level so that the reason an item is ranked at the top can be explained; also, the importance of each user and item review during matching is dynamically adapted. RTM represents users and items based only on their reviews without the need for their identifiers, which improves the generalization ability during product search. dai2023contrastive describe a contrastive learning framework, CoPPS, to enhance user representations for personalized product search. By pre-training a sequence encoder with contrastive sampling and fine-tuning, CoPPS employs multiple data augmentation strategies to improve user modeling.
@@
For real-world applications, zhang2020towards jointly consider semantic matching and personalized matching at JD.com. The authors introduce a deep personalized and semantic retrieval model (DPSR) with a two-tower architecture: a multi-head design of a query tower and an attention-based loss function. Similarly, magnani2022semantic describe a semantic retrieval model with a two-tower architecture in e-commerce search production at Walmart.com. The authors select negative examples for training a large semantic retrieval model and use an approximate metric to evaluate the performance.
@@ 5.4 Ranking strategies in e-commerce search
As we have discussed in Section 5.1.1, learning to rank approaches have been applied to e-commerce search. Unlike web search scenarios based on relevance judgments, e-commerce search has multiple implicit and explicit signals that need to be integrated during the ranking process. In e-commerce search, learning to rank methods need to tackle a series of challenges [karmaker2017application]. Learning to rank methods assume that the scoring of items to be ranked is a parameterized function of multiple features computed based on the given query and the items, whereas parameters are used to control the weights of features. A large number of product and query features, such as brands, rating, categories, etc., are important to obtain useful representations in e-commerce learning to rank models. Also, product popularity related features have been shown to be effective in optimizing ranking results of e-commerce search. karmaker2017application study the performance of multiple learning to rank strategies in e-commerce search, and find that LambdaMART [burges2010ranknet] is able to learn a balance between these two kinds of features. Detailed comparisons are listed in Table 5.1. In addition, user engagement behavior, such as clicks and orders, also plays an important role in optimizing ranking results [karmaker2017application, wu2018turning]. In contrast to web search which only has clicks to judge the relevance, e-commerce search contains four prominent relevance feedback behavior: clicks, cart-adds, orders, and revenue. Thus, various training objectives can be considered in optimizing e-commerce ranking results.
@@
Another main challenge for e-commerce learning to rank methods is the lack of labeled information. In web search, high-quality labeled information obtained by eliciting relevance ratings from human experts or crowdsourcing, makes learning to rank methods effective. However, in the context of e-commerce search it is infeasible to deermine a standard method to obtain ground-truth information. Intuitively, crowd sourcing annotations seem to provide labels for e-commerce learning to rank methods. karmaker2017application and alonso2009relevance study the reliability of the relevance judgements provided by crowd workers for e-commerce queries. The authors find that crowdsourcing fails to provide reliable relevance judgements for e-commerce queries. Thus implicit feedback from e-commerce users is an important source of information that helps reveal the saliency of products for a given query.
@@
wu2018turning divide the ranking optimization problem into two successive stages: click optimization and purchase optimization. We have discussed previous studies on click modeling and purchase modeling in Section 4.1.1 and 4.1.3, respectively. Inspired by these approaches, wu2018turning apply a list-wise learning to rank model [cao2007learning] to optimize clicks by jointly considering item positions and query-level structures; a binary classification approach is applied to predict the purchase behavior.
@@
Online learning to rank approaches have been applied to optimize ranking results in e-commerce search [hu2018reinforcement]. Unlike traditional static learning to rank models, online learning to rank methods optimize the production ranker interactively by exploiting users’ implicit feedback [zoghi-copeland-2015, oosterhuis-balancing-2017, schuth-multileave-2016]. Online learning to rank approaches can be divided into two groups: the first is to learn the best ranking function from a function space [hofmann2016online, hofmann2013balancing, yue2009interactively]; the second group directly learns the best list under some model of user interactions [radlinski2008learning, schuth-multileave-2016, oosterhuis-balancing-2017]. As part of the first group of online learning to rank methods, hu2018reinforcement propose a reinforcement learning method for ranking optimization in e-commerce searching scenarios. The authors formulate the multi-step ranking procedure in e-commerce search as a search session Markov decision process (SSMDP). An algorithm, named deterministic policy gradient with full backup estimation (DPG-FBE), is then proposed for the problem of high reward variance and unbalanced reward distribution of SSMDP. To integrate search results from heterogeneous sources, takanobu2019aggregating introduce a search result aggregation method that formulates a semi-Markov decision process, where a low-level policy is applied to represent items and a high-level policy is used to select rankers. Based on a large-scale real-world dataset, pkddAnwaarRK20 employ a counterfactual risk minimization (CRM) approach to directly optimize the ranking list from the log data.
@@
In recent years, deep neural-based retrieval models [mitra2017introduction, kenter-neural-2017, onal-neural-2018] have been used in the ranking step of e-commerce search. magnani2019neural enhance the deep neural network model using different types of text representation and loss function at Warmart.com. zhang2019neural detail an e-commerce ranking model with interaction features between the query and a graph of products maintaining product interactions. Pre-trained foundation models have been shown to be effective in real-world web search scenarios [zou2021pre, lin-2021-pretrained, chu2022h]. In e-commerce search, wu2022multi apply BERT for product ranking within a multi-task learning framework. The authors use the probability transfer method in the framework to model multiple sequential engagement behaviors. By integrating semantic matching features output by the domain-specific BERT, the authors confirm the effectiveness of their proposed approach on real-world e-commerce search data.
@@ 5.5.1 Multi-modal e-commerce search
Along with the rise of deep neural networks, multi-modal search has increasingly received attention [zhang2013attribute, mao2014deep, yang2017visual, liu2018attentive, balaneshin2018deep, guo2018multi, zhang2018visual, qu2021dynamic, wei2021universal, tan2022bit]. The key of multi-modal search is to find an effective mapping mechanism to project data from different modalities into a common latent space. Multi-modal search approaches can be classified into hashing-aware models and semantic-aware models. The former type of methods map various modalities in the original space to a Hamming space using hash functions [cao2017transitive, luo2018scalable, zhu2020deep, tan2022bit], whereas the latter ones project the multi-modal data into a low-dimensional space by learning a mapping function [wang2016effective, laenen2018web, qu2021dynamic, wang2022siamese].
@@
As we discussed in previous sections, most approaches to e-commerce search focus on the textual matching problem. With the rise in online photos and openly available image datasets this is changing. yang2017visual propose a novel end-to-end approach for scalable visual search infrastructure at Ebay. Similar platforms can be found at Alibaba [zhang2018visual] and JD.com [li2018design, wang2020metasearch]. Early studies into multi-modal e-commerce search have mostly focused on the fashion category [yang2017visual, zhang2018visual, li2018design]. However, wang2020metasearch, sigir21DaganGN21, liu2022pretraining confirmed that multi-modal search is widespread across many e-commerce categories, especially categories that involve aspects that are harder to express verbally, but can naturally be captured visually, such as style, type, and pattern.
@@
In multi-modal search, multiple modalities, such as text and images, can be found in both queries and products. Most multi-modal search solutions optimize a function to project multi-modal data into a low-dimensional space after unified representation learning. laenen2018web present a multi-modal search paradigm for e-commerce search that results in an improved shopping experience. The authors reason with both images and languages through a common embedding space. guo2018multi formulate the e-commerce personalized search problem based on the relevance between images and text with respect to the query and the user preferences from both textual and visual modalities. A transition-based product search method has been proposed, where the multi-modal feature space is initialized based on the textual and visual features of products. An interpretable multi-modal e-commerce retrieval framework has been proposed for fashion products [liao2018interpretable]; the authors bridge the gap between deep features and meaningful fashion concepts. They propose a hierarchical similarity function to accurately characterize the semantic affinities among fashion items.
@@
sigir21DaganGN21 highlight various differences between visual and textual search, which can be summarized in the following challenges in multi-modal e-commerce search:
@@
  • Large-scale sparse data. Large multimedia corpora make scalability and efficiency key requirements for e-commerce search. Visual queries are more specific than text queries, which results in a smaller number of retrieved results and sparse coverage of categories.
  • Limited user engagement. In multi-modal e-commerce search, user engagement behavior, e.g., clicks, dwell time, purchases, etc., is substantially sparser than in textual search.
  • Ambiguous user intents. Two different main use cases exist in multi-modal e-commerce search: target finding and decision making [sigir21DaganGN21]. The two use cases reflect the navigational intent and informational intent, respectively.
Additionally, annotating new product images and retraining a new feature representation model on large-scale data is expensive. To address this problem, few-shot multi-modal product search has been proposed to update the feature representation and index model with few-shot data and employ a fast learning strategy for new categories. wang2020metasearch describe a framework for few-shot incremental search via meta-learning, with a multi-pooling feature extractor to extract discriminative multi-modal features. Based on pre-trained language models, liu2022pretraining detail an effective contrastive learning framework to learn representations of multi-modal search sessions based on multi-view heterogeneous graph networks. liu2023multimodal employ BERT with a self-distillation framework for product understanding by integrating visual and textual information.
Cross-modal search is receiving more and more attention [qu2021dynamic, wang2022siamese, tan2022bit]. Unlike multi-modal search, cross-modal search aims to solve the discrepancy problem between different modalities in search sessions [wang2022siamese]. Although much progress has been made in bridging multiple modalities, it still remains challenging because of the difficult intra-modal reasoning and cross-modal alignment [qu2021dynamic]. As complicated relations exist among products in e-commerce, it is more difficult to explore these multi-hop interactions in cross-modal product search scenarios. Moreover, the extremely high computational costs brought by multi-modal input limits its usefulness in large-scale cross-modal retrieval in e-commerce applications.